# Optimal Learners for Realizable Regression: PAC Learning and Online Learning

### v1

@@ Abstract
Previous work had established the sufficiency of finiteness of the fat shattering dimension for PAC learnability and the necessity of finiteness of the scaled Natarajan dimension, but little progress had been made towards a more complete characterization since the work of Simon 1997 (SICOMP ’97). To this end, we first introduce a minimax instance optimal learner for realizable regression and propose a novel dimension that both qualitatively and quantitatively characterizes which classes of real-valued predictors are learnable. We then identify a combinatorial dimension related to the Graph dimension that characterizes ERM learnability in the realizable setting. Finally, we establish a necessary condition for learnability based on a combinatorial dimension related to the DS dimension, and conjecture that it may also be sufficient in this context.
@@ 1 Introduction
Perhaps surprisingly, in the fundamental realizable Probably Approximately Correct (PAC) setting [Val84] and the realizable online setting [Lit88, DG22], we do not know of any characterizing dimension or optimal learners for the regression task. This comes in sharp contrast with binary and multiclass classification, both in the offline and the online settings, where the situation is much more clear [Lit88, HLW94, BEHW89, Han16, DSS14, ${}^{+}$22, Nat89, DSBDSS15, RST23]. Our goal in this work is to make progress regarding the following important question:
@@
We next shift our attention to the classical setting of online learning where the learner interacts with the adversary over a sequence of $T$ rounds: in every round the adversary presents an example $x_{t}\in\mathcal{X}$, the learner predicts a label $\widehat{y}_{t}$ and then the adversary reveals the correct label $y^{\star}_{t}.$ In this context, realizability means that there always exists some function $h_{t}\in\mathcal{H}$ that perfectly explains the examples and the labels that the adversary has chosen. In the agnostic setting, the goal of the learner is to compete with the performance of the best function in $\mathcal{H}$, i.e., achieve small regret. This setting was introduced by [Lit88] in the context of binary classification, where they also characterized learnability in the realizable setting and provided an optimal algorithm. Later, [BDPSS09] provided an almost optimal algorithm in the agnostic setting, which suffered from an additional $\log T$ factor in its regret bound. This extra factor was later shaved by [${}^{+}$21]. The study of multiclass online classification was initiated by [DSBDSS15], who provided an optimal algorithm for the realizable setting and an algorithm that is suboptimal by a factor of $\log k\cdot\log T$ in the agnostic setting, where $k$ is the total number of labels. Recently, [RST23] shaved off the $\log k$ factor. The problem of online regression differs significantly from that of online classification, since the loss function is not binary. The agnostic online regression setting has received a lot of attention and there is a series of works that provides optimal minimax guarantees [RST10, RS14, RST15b, RST15a, BDR21].
@@ 1.1 Related Work
Our work makes progress towards the characterization of realizable regression both in the offline and online settings. Similar results are known for the more studied settings of binary and multiclass (PAC/online) classification [Lit88, HLW94, BEHW89, Han16, DSS14, ${}^{+}$22, Nat89, DSBDSS15, RBR09, BLM20, RST23, ${}^{+}$21, HKLM22, KKMV23, ${}^{+}$23]. The fundamental works of [BLW94, Sim97, ABDCBH97, BL98] study the question of PAC learnability for the regression task. However, none of them characterizes learnability in the realizable setting. [Sim97] showed that finiteness of scaled Natarajan dimension is necessary for realizable PAC regression. [BL98] employed the one-inclusion graph (OIG) algorithm to get a real-valued predictor whose expected error is upper bounded by the $V_{\gamma}$-dimension, whose finiteness is sufficient but not necessary for realizable learnability in this setting. We refer to [KS23] for details about this dimension which was introduced in [ABDCBH97]. [ABDCBH97] showed that finiteness of the fat shattering dimension at all scales is equivalent to PAC learnability in the agnostic setting. Nevertheless, this does not hold in the realizable case as Example 1 demonstrates. Similarly, the work of [BLW94] shows that fat shattering dimension characterizes the regression task when the labels are corrupted by noise. Recently, the concurrent and independent work of [AACSZ23] provides high-probability bounds for the one-inclusion graph algorithm in the realizable PAC regression setting using the $V_{\gamma}$-dimension. We underline that this dimension does not characterize realizable regression. For more general losses, see [Men02, BBM05].
@@
Our characterization (cf. Theorem 2) for the offline setting via the OIG algorithm [HLW94] is further motivated by the work of [MHS22], where they propose a dimension of similar flavor for adversarially robust PAC learnability. We mention that recently the OIG algorithm has received a lot of attention from the statistical learning theory community [AACSZ22, ${}^{+}$21, KVK22, SMB22, AHHM22, CP22]. Finally, for a small sample of works that deal with offline and online regression problems, see [BDGR22, Gol21, AH22, HKS19, She17, MKFI22] and the references therein.
@@ 2.1 $\gamma$-Fat Shattering Dimension
We remark that this example already shows that the PAC learnability landscape of regression is quite different from that of multiclass classification, where agnostic learning and realizable learning are characterized by the same dimension [${}^{+}$22].
@@ 2.2 $\gamma$-Natarajan Dimension
The $\gamma$-Natarajan dimension was introduced by [Sim97] and is inspired by the Natarajan dimension [Nat89], which has been used to derive bounds in the multiclass classification setting.
@@ 2.4 $\gamma$-One-Inclusion Graph Dimension
In this section, we provide a minimax optimal learner for realizable PAC regression. We first review a fundamental construction which is a crucial ingredient in the design of our learner, namely the one-inclusion (hyper)graph (OIG) algorithm $\mathbb{A}^{\mathrm{OIG}}$ for a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ [HLW94, RBR09, DSS14, ${}^{+}$22]. This algorithm gets as input a training set $(x_{1},y_{1}),...,(x_{n},y_{n})$ realizable by $\mathcal{H}$ and an additional example $x$. The goal is to predict the label of $x$. Let $S=\{x_{1},...,x_{n},x\}$. The idea is to construct the one-inclusion graph $G_{\mathcal{H}|_{S}}^{\mathrm{OIG}}$ induced by the pair $(S,\mathcal{H})$. The node set $V$ of this graph corresponds to the set $\mathcal{H}|_{S}$ (projection of $\mathcal{H}$ to $S$) and, so, $V\subseteq\mathcal{Y}^{[n+1]}$. For the binary classification case, two vertices are connected with an edge if they differ in exactly one element $x$ of the $n+1$ points in $S$. For the case where $\mathcal{Y}$ is discrete and $|\mathcal{Y}|>2$, the hyperedge is generalized accordingly.
Definition 7 (One-Inclusion Hypergraph [HLW94, RBR09, ${}^{+}$22]).
Consider the set $[n]$ and a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{[n]}$. We define a graph $G^{\mathrm{OIG}}_{\mathcal{H}}=(V,E)$ such that $V=\mathcal{H}$. Consider a direction $i\in[n]$ and a mapping $f:[n]\setminus\{i\}\to\mathcal{Y}$. We introduce the hyperedge $e_{i,f}=\{h\in V:h(j)=f(j),~{}\forall i,j\in[n]^{2},i\neq j\}$. We define the edge set of $G^{\mathrm{OIG}}_{\mathcal{H}}$ to be the collection
@@
In words, it is the largest $n\in\mathbb{N}$ (potentially $\infty$) such that there exists an (unlabeled) sequence $S$ of length $n$ with the property that no matter how one orients some finite subgraph of the one-inclusion graph, there is always some vertex for which at least $1/3$ of its coordinates are $\gamma$-different from the labels of the edges that are attached to this vertex. We remark that the hypothesis class in Example 1 has $\gamma$-OIG dimension equal to $O(1)$. We also mention that the above dimension satisfies the “finite character” property and the remaining criteria that dimensions should satisfy according to [${}^{+}$19] (see Appendix F). As our main result in this section, we show that any class $\mathcal{H}$ is learnable if and only if this dimension is finite and we design an (almost) optimal learner for it.
@@ 2.5 $\gamma$-DS Dimension
So far we have identified a dimension (cf. Definition 9) that characterizes the PAC learnability of realizable regression. However, it might not be easy to calculate it in some settings. Our goal in this section is to introduce a relaxation of this definition which we conjecture that also characterizes learnability in this context. This new dimension is inspired by the DS dimension, a combinatorial dimension defined by Daniely and Shalev-Shwartz in [DSS14]. In a recent breakthrough result, [${}^{+}$22] showed that the DS dimension characterizes multiclass learnability (with a possibly unbounded number of labels and the 0-1 loss). We introduce a scaled version of the DS dimension. To this end, we first define the notion of a scaled pseudo-cube.
@@
Let $\gamma\in[0,1]$. A class $\mathcal{H}\subseteq[0,1]^{d}$ is called a $\gamma$-pseudo-cube of dimension $d$ if it is non-empty, finite and, for any $f\in\mathcal{H}$ and direction $i\in[d]$, the hyper-edge $e_{i,f}=\{g\in\mathcal{H}:g(j)=f(j)~{}~{}\forall i,j\in[d]^{2},i\neq j\}$ satisfies $|e_{i,f}|>1$ and $\ell(g_{1}(i),g_{2}(i))>\gamma$ for any $g_{1},g_{2}\in e_{i,f},g_{1}\neq g_{2}$.
Pseudo-cubes can be seen as a relaxation of the notion of a Boolean cube (which should be intuitively related with the Natarajan dimension) and were a crucial tool in the proof of [${}^{+}$22]. In our setting, scaled pseudo-cubes will give us the following combinatorial dimension.
@@
The proof is postponed to Appendix D. We believe that finiteness of $\mathbb{D}^{\mathrm{DS}}_{\gamma}(\mathcal{H})$ is also a sufficient condition for realizable regression. However, the approach of [${}^{+}$22] that establishes a similar result in the setting of multiclass classification does not extend trivially to the regression setting.
@@ 4 Conclusion
In this work, we developed optimal learners for realizable regression in PAC learning and online learning. Moreover, we identified combinatorial dimensions that characterize learnability in these settings. We believe that our work can lead to simplified characterizations for these problems.

### v3

@@ Abstract
Previous work had established the sufficiency of finiteness of the fat shattering dimension for PAC learnability and the necessity of finiteness of the scaled Natarajan dimension, but little progress had been made towards a more complete characterization since the work of Simon (SICOMP ’97). To this end, we first introduce a minimax instance optimal learner for realizable regression and propose a novel dimension that both qualitatively and quantitatively characterizes which classes of real-valued predictors are learnable. We then identify a combinatorial dimension related to the Graph dimension that characterizes ERM learnability in the realizable setting. Finally, we establish a necessary condition for learnability based on a combinatorial dimension related to the DS dimension, and conjecture that it may also be sufficient in this context.
@@ 1 Introduction
Perhaps surprisingly, in the fundamental realizable Probably Approximately Correct (PAC) setting [Val84] and the realizable online setting [Lit88, DG22], we do not know of any characterizing dimension or optimal learners for the regression task. This comes in sharp contrast with binary and multiclass classification, both in the offline and the online settings, where the situation is much more clear [Lit88, HLW94, BEHW89, Han16, DSS14, BCD+22, Nat89, DSBDSS15, RST23]. Our goal in this work is to make progress regarding the following important question:
@@
We next shift our attention to the classical setting of online learning where the learner interacts with the adversary over a sequence of $T$ rounds: in every round the adversary presents an example $x_{t}\in\mathcal{X}$, the learner predicts a label $\widehat{y}_{t}$ and then the adversary reveals the correct label $y^{\star}_{t}.$ In this context, realizability means that there always exists some function $h_{t}\in\mathcal{H}$ that perfectly explains the examples and the labels that the adversary has chosen. In the agnostic setting, the goal of the learner is to compete with the performance of the best function in $\mathcal{H}$, i.e., achieve small regret. This setting was introduced by [Lit88] in the context of binary classification, where they also characterized learnability in the realizable setting and provided an optimal algorithm. Later, [BDPSS09] provided an almost optimal algorithm in the agnostic setting, which suffered from an additional $\log T$ factor in its regret bound. This extra factor was later shaved by [ABED+21]. The study of multiclass online classification was initiated by [DSBDSS15], who provided an optimal algorithm for the realizable setting and an algorithm that is suboptimal by a factor of $\log k\cdot\log T$ in the agnostic setting, where $k$ is the total number of labels. Recently, [RST23] shaved off the $\log k$ factor. The problem of online regression differs significantly from that of online classification, since the loss function is not binary. The agnostic online regression setting has received a lot of attention and there is a series of works that provides optimal minimax guarantees [RST10, RS14, RST15b, RST15a, BDR21].
@@ 1.1 Related Work
Our work makes progress towards the characterization of realizable regression both in the offline and online settings. Similar results are known for the more studied settings of binary and multiclass (PAC/online) classification [Lit88, HLW94, BEHW89, Han16, DSS14, BCD+22, Nat89, DSBDSS15, RBR09, BLM20, RST23, ABED+21, HKLM22, KKMV23, BGH+23]. The fundamental works of [BLW94, Sim97, ABDCBH97, BL98] study the question of PAC learnability for the regression task. However, none of them characterizes learnability in the realizable setting. [Sim97] showed that finiteness of scaled Natarajan dimension is necessary for realizable PAC regression. [BL98] employed the one-inclusion graph (OIG) algorithm to get a real-valued predictor whose expected error is upper bounded by the $V_{\gamma}$-dimension, whose finiteness is sufficient but not necessary for realizable learnability in this setting. We refer to [KS23] for details about this dimension which was introduced in [ABDCBH97]. [ABDCBH97] showed that finiteness of the fat shattering dimension at all scales is equivalent to PAC learnability in the agnostic setting. Nevertheless, this does not hold in the realizable case as Example 1 demonstrates. Similarly, the work of [BLW94] shows that fat shattering dimension characterizes the regression task when the labels are corrupted by noise. Recently, the concurrent and independent work of [AACSZ23] provides high-probability bounds for the one-inclusion graph algorithm in the realizable PAC regression setting using the $V_{\gamma}$-dimension. We underline that this dimension does not characterize realizable regression. For more general losses, see [Men02, BBM05].
@@
Our characterization (cf. Theorem 2) for the offline setting via the OIG algorithm [HLW94] is further motivated by the work of [MHS22], where they propose a dimension of similar flavor for adversarially robust PAC learnability. We mention that recently the OIG algorithm has received a lot of attention from the statistical learning theory community [AACSZ22, BHM+21, KVK22, SMB22, AHHM22, CP22]. Finally, for a small sample of works that deal with offline and online regression problems, see [BDGR22, Gol21, AH22, HKS19, She17, MKFI22] and the references therein.
@@ 2.1 $\gamma$-Fat Shattering Dimension
We remark that this example already shows that the PAC learnability landscape of regression is quite different from that of multiclass classification, where agnostic learning and realizable learning are characterized by the same dimension [BCD+22].
To summarize this subsection, the fat-shattering dimension is a natural way to quantify how well the function class can interpolate (with gap $\gamma$) some fixed function. Crucially, this interpolation contains only inequalities (see Definition 4) and hence (at least intuitively) cannot be tight for the realizable setting, where there exists some function that exactly labels the features. Example 1 gives a natural example of a class with infinite fat-shattering dimension that can, nevertheless, be learned with a single sample in the realizable setting.
@@ 2.2 $\gamma$-Natarajan Dimension
The $\gamma$-Natarajan dimension was introduced by [Sim97] and is inspired by the Natarajan dimension [Nat89], which has been used to derive bounds in the multiclass classification setting. Before explaining the $\gamma$-Natarajan dimension, let us begin by reminding to the reader the standard Natarajan dimension. We say that a set $S=\{x_{1},...,x_{n}\}$ of size $n$ is Natarajan-shattered by a concept class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ if there exist two functions $f,g:S\to\mathcal{Y}$ so that $f(x_{i})\neq g(x_{i})$ for all $i\in[n]$, and for all $b\in\{0,1\}^{n}$ there exists $h\in\mathcal{H}$ such that $h(x_{i})=f(x_{i})$ if $b_{i}=1$ and $h(x_{i})=g(x_{i})$ if $b_{i}=0$. Note that here we have equalities instead of inequalities (recall the fat-shattering case Definition 4).
From a geometric perspective (see [BCD+22]), this means that the space $\mathcal{H}$ projected on the set $S$ contains the set $\{f(x_{1}),g(x_{1})\}\times...\times\{f(x_{n}),g(x_{n})\}$. This set is "isomorphic" to the Boolean hypercube of size $n$ by mapping $f(x_{i})$ to 1 and $g(x_{i})$ to 0 for all $i\in[n]$. This means that the Natarajan dimension is essentially the size of the largest Boolean cube contained in $\mathcal{H}$.
[Sim97] defines the scaled analogue of the above dimension as follows.
@@ 2.4 $\gamma$-One-Inclusion Graph Dimension
In this section, we provide a minimax optimal learner for realizable PAC regression. We first review a fundamental construction which is a crucial ingredient in the design of our learner, namely the one-inclusion (hyper)graph (OIG) algorithm $\mathbb{A}^{\mathrm{OIG}}$ for a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ [HLW94, RBR09, DSS14, BCD+22]. This algorithm gets as input a training set $(x_{1},y_{1}),...,(x_{n},y_{n})$ realizable by $\mathcal{H}$ and an additional example $x$. The goal is to predict the label of $x$. Let $S=\{x_{1},...,x_{n},x\}$. The idea is to construct the one-inclusion graph $G_{\mathcal{H}|_{S}}^{\mathrm{OIG}}$ induced by the pair $(S,\mathcal{H})$. The node set $V$ of this graph corresponds to the set $\mathcal{H}|_{S}$ (projection of $\mathcal{H}$ to $S$) and, so, $V\subseteq\mathcal{Y}^{[n+1]}$. For the binary classification case, two vertices are connected with an edge if they differ in exactly one element $x$ of the $n+1$ points in $S$. For the case where $\mathcal{Y}$ is discrete and $|\mathcal{Y}|>2$, the hyperedge is generalized accordingly.
Definition 7 (One-Inclusion Hypergraph [HLW94, RBR09, BCD+22]).
Consider the set $[n]$ and a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{[n]}$. We define a graph $G^{\mathrm{OIG}}_{\mathcal{H}}=(V,E)$ such that $V=\mathcal{H}$. Consider a direction $i\in[n]$ and a mapping $f:[n]\setminus\{i\}\to\mathcal{Y}$. We introduce the hyperedge $e_{i,f}=\{h\in V:h(j)=f(j),~{}\forall j\in[n]\setminus\{i\}\}$. We define the edge set of $G^{\mathrm{OIG}}_{\mathcal{H}}$ to be the collection
@@
In words, it is the largest $n\in\mathbb{N}$ (potentially $\infty$) such that there exists an (unlabeled) sequence $S$ of length $n$ with the property that no matter how one orients some finite subgraph of the one-inclusion graph, there is always some vertex for which at least $1/3$ of its coordinates are $\gamma$-different from the labels of the edges that are attached to this vertex. We remark that the hypothesis class in Example 1 has $\gamma$-OIG dimension equal to $O(1)$. We also mention that the above dimension satisfies the “finite character” property and the remaining criteria that dimensions should satisfy according to [BDHM+19] (see Appendix F). As our main result in this section, we show that any class $\mathcal{H}$ is learnable if and only if this dimension is finite and we design an (almost) optimal learner for it.
@@
Before proceeding to the next section, one could naturally ask whether there is a natural property of the concept class that implies finiteness of the scaled OIG dimension. The work of [Men02] provides a sufficient and natural condition that implies finiteness of our complexity measure. In particular, Mendelson shows that classes that contain functions with bounded oscillation (as defined in [Men02]) have finite fat-shattering dimension. This implies that the class is learnable in the agnostic setting and hence is also learnable in the realizable setting. As a result, the OIG-based dimension is also finite. So, bounded oscillations are a general property that guarantees that the finiteness of OIG-based dimension and fat-shattering dimension coincide. We also mention that deriving bounds for the OIG-dimension for interesting families of functions is an important yet non-trivial question.
A finite fat-shattering dimension implies a finite OIG dimension.
Let $\mathcal{F}\subseteq[0,1]^{\mathcal{X}}$ be a function class with finite $\gamma$-fat shattering dimension for any $\gamma>0$. We show that $\mathbb{D}^{\mathrm{OIG}}_{\gamma}(\mathcal{F})$ is upper bounded (up to constants and log factors) by $\mathbb{D}^{\mathrm{fat}}_{c\gamma}(\mathcal{F})$, for some $c>0$, where the OIG dimension is defined with respect to the $\ell_{1}$ loss. Note that the opposite direction does not hold. Example 1 exhibits a function class with an infinite fat-shattering dimension that can be learned with a single example, and as a result, the OIG dimension has to be finite. On the one hand, we have an upper bound on the sample complexity of $O\left(\frac{1}{\epsilon}\left(\mathbb{D}^{\mathrm{fat}}_{\tilde{c}\varepsilon}(\mathcal{F})\log^{2}\frac{1}{\epsilon}+\log\frac{1}{\delta}\right)\right)$, for any $\varepsilon,\delta\in(0,1)$. See sections 19.6 and 20.4 about the restricted model in [AB99]. On the other hand, we prove in Lemma 6 a lower bound on the sample complexity of $\Omega\left(\frac{\mathbb{D}^{\mathrm{OIG}}_{2\varepsilon}(\mathcal{F})}{\epsilon}\right)$, for any $\varepsilon\in(0,1)$, and so $\mathbb{D}^{\mathrm{OIG}}_{\gamma}(\mathcal{F})$ is upper bounded by $\mathbb{D}^{\mathrm{fat}}_{c\gamma}(\mathcal{F})$ up to constants and log factors.
@@ 2.5 $\gamma$-DS Dimension
So far we have identified a dimension (cf. Definition 9) that characterizes the PAC learnability of realizable regression. However, it might not be easy to calculate it in some settings. Our goal in this section is to introduce a relaxation of this definition which we conjecture that also characterizes learnability in this context. This new dimension is inspired by the DS dimension, a combinatorial dimension defined by Daniely and Shalev-Shwartz in [DSS14]. In a recent breakthrough result, [BCD+22] showed that the DS dimension characterizes multiclass learnability (with a possibly unbounded number of labels and the 0-1 loss). We introduce a scaled version of the DS dimension. To this end, we first define the notion of a scaled pseudo-cube.
@@
Let $\gamma\in[0,1]$. A class $\mathcal{H}\subseteq[0,1]^{d}$ is called a $\gamma$-pseudo-cube of dimension $d$ if it is non-empty, finite and, for any $f\in\mathcal{H}$ and direction $i\in[d]$, the hyper-edge $e_{i,f}=\{g\in\mathcal{H}:g(j)=f(j)~{}~{}\forall j\in[d],i\neq j\}$ satisfies $|e_{i,f}|>1$ and $\ell(g_{1}(i),g_{2}(i))>\gamma$ for any $g_{1},g_{2}\in e_{i,f},g_{1}\neq g_{2}$.
Pseudo-cubes can be seen as a relaxation of the notion of a Boolean cube (which should be intuitively related with the Natarajan dimension) and were a crucial tool in the proof of [BCD+22]. In our setting, scaled pseudo-cubes will give us the following combinatorial dimension.
@@
The proof is postponed to Appendix D. We believe that finiteness of $\mathbb{D}^{\mathrm{DS}}_{\gamma}(\mathcal{H})$ is also a sufficient condition for realizable regression. However, the approach of [BCD+22] that establishes a similar result in the setting of multiclass classification does not extend trivially to the regression setting.
@@ 4 Conclusion
In this work, we developed optimal learners for realizable regression in PAC learning and online learning. Moreover, we identified combinatorial dimensions that characterize learnability in these settings. We hope that our work can lead to simplified characterizations for these problems. We believe that the main limitation of our work is that the OIG-based dimension we propose is more complicated than the dimensions that have been proposed in the past, like the fat-shattering dimension (which, as we explain, does not characterize learnability in the realizable regression setting). Nevertheless, despite its complexity, this is the first dimension that characterizes learnability in the realizable regression setting. More to that, our work leaves as an important next step to prove (or disprove) the conjecture that the (combinatorial and simpler) $\gamma$-DS dimension is qualitatively equivalent to the $\gamma$-OIG dimension. Another future direction, that is not directly related to this conjecture, is to better understand the gap between the fat-shattering dimension and the OIG-based dimension.
Acknowledgements
Amin Karbasi acknowledges funding in direct support of this work from NSF (IIS-1845032), ONR (N00014- 19-1-2406), and the AI Institute for Learning-Enabled Optimization at Scale (TILOS). Grigoris Velegkas is supported by TILOS, the Onassis Foundation, and the Bodossaki Foundation. This work was done in part while some of the authors were visiting Archimedes AI Research Center.